# Conditional Density Estimations from Privacy-Protected Data

### v3

@@ 5 Discussion
Experimental setup. We employed neural spline flows (Durkan et al.,, 2019) as the conditional density estimator, consisting of 8 layers. Each layer was constructed using two residual blocks with 50 units and ReLU activation function, with 10 bins in each monotonic piecewise rational-quadratic transforms and the tail bound was set to 5.
In the training process, the number of samples simulated in each round is $N=1000$ and there are $R=10$ rounds in total. In each round of training, we randomly select 5% of the newly generated samples as validation data. According to the early stop criterion proposed by Papamakarios et al., (2019), we stop training if the value of loss on validation data does not increase after 20 epochs in a single round. For stochastic gradient descent optimizer, we choose the Adam (Kingma and Ba,, 2014) with the batchsize of 100, the learning rate of $5\times 10^{-4}$ and the weight decay is $10^{-4}$.
S3.1 SIR model
S3.1.1 Detailed results on synthetic data
In our experiments, we use the Gillespie algorithm (Gillespie,, 1977) to simulate the whole process over a duration of $T=160$ time units, and recorded the populations at intervals of 1 time units. The prior distribution of $\beta$ is set to $\mathcal{N}(\log 0.4,0.5)$ and prior distribution of $\gamma$ is set to $\mathcal{N}(\log 0.125,0.2)$. The value of $K$ is configured as 1000000, while $N$ is set to 1000 for the number of observations.
To publish the privatized data about the infection process, we select the infectious group $I(t)$ at $L=10$ evenly-spaced points in time $[0,T]$, with the privacy parameters set to $n=1000$ and $m=1000$, which satisfied $\epsilon$-DP with $\epsilon=10$. The ground truth parameters is
@@
Figure S2 illustrates the convergence of the approximate posterior by SPPE and SPLE in each round, and compares our results with the SMC-ABC method, where we performed simulations up to $5\times 10^{5}$ times for the SMC-ABC method to generate the near exact ‘True’ posterior. Figure S1 depicts the results of the SMC-ABC method under the same performance metrics. Our method, after 10 rounds or $10^{4}$ simulations, achieved similar performance as the SMC-ABC method with approximately $10^{5}$ simulations.
Table S1: Computational Cost to achieve MMD $<$ 0.1 in the SIR Model (Mean $\pm$ Standard Deviation)
Method Simulation Time Network Training Time Total Number of Simulations
@@
Figure S1: Approximation accuracy by SMC-ABC on the SIR model against number of simulations.
Figure S2: Detailed convergence of sequential posterior estimations given DP-protected infection trajectory under the SIR model. Each round entails $N=1000$ simulations. Orange: SPPE; red: SPLE; grey: SMC-ABC.
Finally, we investigate the marginal posterior distributions $\pi(\beta\mid s_{\mathrm{dp}})$ and $\pi(\gamma\mid s_{\mathrm{dp}})$ in Figure S3, and conclude that SPPE and SPLE perform similarly well. In this example, the posterior approximated by SPPE is slightly more concentrated than those from SMC-ABC and SPLE.
Figure S3: Marginal posterior histograms of $\beta,\gamma$, and $\beta/\gamma$ in SIR model on synthetic data. Grey: SMC-ABC; orange: SPPE; red: SPLE. The vertical lines indicate true data generating parameters, set to mimic a measles outbreak.
S3.1.2 Inference on real disease outbreaks
We applied our inference methods (SPPE and SPLE) to analyze several real infectious disease outbreaks, namely influenza, Ebola, and Covid-19.
influenza outbreak. We utilized the dataset from a boarding school, obtained from https://search.r-project.org/CRAN/refmans/epimdr/html/flu.html. The total population in the school was $K=763$. We considered a daily time interval, and the observed private statistic $s_{\mathrm{dp}}^{o}$ is
@@
Ebola outbreak in West Africa, 2014. We analyzed datasets from three regions: a) Guinea, b) Liberia, and c) Sierra Leone. The dataset source is from https://usafacts.org/visualizations/coronavirus-covid-19-spread-map/state/nevada/county/clark-county/. We assumed a potential contact individuals of $K=100,000$. We selected 9 equally spaced time intervals of 120 days starting from 03/31/2014. The resulting observed private statistic $s_{\mathrm{dp}}^{o}$ is as follows
@@
Covid-19. We examined the Covid-19 dataset for Clark County, Nevada. See https://usafacts.org/visualizations/coronavirus-covid-19-spread-map/state/nevada/county/clark-county/. We assumed a potential contact population of $K=100,000$. We selected 9 equally spaced time intervals of 24 days, starting from 09/07/2020. To obtain the number of currently infected individuals, we calculated the difference in the total confirmed cases with a time interval of 14 days from the original dataset. The resulting observed private statistic $s_{\mathrm{dp}}^{o}$ is:
@@
S3.2 Bayesian linear regression
Data generating parameters. Following the parameters settings in Ju et al., (2022), we model the predictors with $x_{0,i}\sim\mathcal{N}_{p}(m,\Sigma)$, where $\Sigma=I_{n}$ and $m=(0.9,\ -1.17)$. The outcomes $y$ satisfy $y\mid x_{0}\sim\mathcal{N}_{n}(x\beta,\sigma^{2}I_{n})$ where $\sigma^{2}=2$, and the prior for $\beta$ is independent normal $\mathcal{N}(0,1)$. We set the privacy loss budget $\epsilon=10$ and the number of subjects $n=100$. The ground truth parameters are denoted as
@@
Experimental results. In Figure S4, we present a comparison of the performance of the SPPE and SPLE methods across different metrics as the number of simulation rounds increases. The SPLE method demonstrates a faster convergence in this task. Figure S5 displays the posterior after 10 rounds, where both the SPPE and SPLE methods achieve results that are close to the near exact posterior obtained by the SMC-ABC method.
Table S2: Estimated Posterior Mean and 95% Credible Intervals for the Linear Regression Example with $\epsilon=3$.
@@
SMC-ABC -1.03 (-2.60, 0.53) -1.25 (-2.77, 0.49) 1.19 (-0.33, 2.64)
DA-MCMC -0.99 (-2.71, 0.73) -1.37 (-2.78, 0.43) 1.11 (-0.33, 2.54)
SPPE -1.01 (-2.59, 0.72) -1.28 (-2.83, 0.39) 1.22 (-0.30, 2.72)
SPLE -1.03 (-2.73, 0.67) -1.22 (-2.81, 0.66) 1.23 (-0.31, 2.70)
Gibbs-SS -0.43 (-2.36, 1.54) -0.20 (-1.90, 1.71) 0.25 (-1.51, 2.10)
RNPE -1.34 (-3.48, 1.40) -1.21 (-3.31, 1.78) 1.02 (-1.59, 3.38)
Figure S4: Approximation accuracy by SPPE (orange) and SPLE (red) on the Bayesian linear regression model against number of rounds, the error bars represent the mean with the upper and lower quartiles.
Figure S5: Posterior comparison on the Bayesian linear regression model. Grey: SMC-ABC; orange: SPPE; red: SPLE. The vertical lines and black dots indicate true data generating parameters.
S3.3 Naïve Bayes log-linear model
@@
Model description. The naïve Bayes log-linear model is a commonly used approach for modeling categorical data (Karwa et al.,, 2015). The input feature-vector, denoted as $x=(x_{1},\ldots,x_{K})$, consists of $K$ features, each taking values in the range $\{1,2,\ldots,J_{k}\}$. The output class, denoted as $y$, represents the target category and takes values in $\{1,2,\ldots,I\}$. The model assumes that the conditional probability of the input given the output, denoted as $p(x\mid y)$, can be factorized as the product of individual feature probabilities: $p(x\mid y)=\prod_{k=1}^{K}p(x_{k}\mid y)$. The model parameters are $p_{ij}^{k}$, which represents the probability $p(x_{k}=j\mid y=i)$, with prior $(p_{i,1}^{k},\cdots,p_{i,J_{k}}^{k})\sim\mathrm{Dirichlet}(\alpha_{i,1}^{k},\cdots,\alpha_{i,J_{k}}^{k})$ for all $i$ and $k$; and $p_{i}=p(y=i)$, with prior $(p_{1},\cdots,p_{I})\sim\mathrm{Dirichlet}(\alpha_{1},\cdots,\alpha_{I})$.
@@
We assume that $(n_{i,1}^{k},\cdots,n_{i,J_{k}}^{k})\sim\mathrm{Multinomial}(n_{i};p_{i,1}^{k},\cdots,p_{i,J_{k}}^{k})$ for all $i$ and $k$ and $(n_{1},\cdots,n_{I})\sim\mathrm{Multinomial}(n;p_{1},\cdots,p_{I})$. Here $n_{i,j}^{k}$ represents the counts #$\#(y=i,x_{k}=j)$. One sufficient statistics of the model are the proportion of the counts $r_{i,j}^{k}:=\frac{1}{n}n_{i,j}^{k}$, where $n=\sum_{i=1}^{I}\sum_{j=1}^{J_{k}}n_{i,j}^{k}$ for all $k$. To protect the privacy of the dataset, Laplace noise $e_{i,j}^{k}$ is added to the proportion of the counts, resulting in the privatized proportion $m_{i,j}^{k}=r_{i,j}^{k}+e_{i,j}^{k}$. When $e_{i,j}^{k}\sim\mathrm{Laplace}(0,\frac{2K}{n\epsilon})$, the released private statistic $\{m_{i,j}^{k}\}_{i,j,k}$ satisfied $\epsilon$-DP.
@@
Experimental results. Figure S6 illustrates the performance of the SPPE and SPLE methods across four different metrics, and Figure S7 shows the marginal posterior histograms after 10 rounds, our methods stabilize in performance after round 3.
Figure S6: Approximation accuracy by SPPE (orange) and SPLE (red) on the log-linear model against number of rounds, the error bars represent the mean with the upper and lower quartiles.
Figure S7: Marginal posterior histograms of the log-linear model. Grey: SMC-ABC; orange: SPPE; red: SPLE. The vertical lines indicate true data generating parameters.
S4 Statement on Computing Resources
Our numerical experiments were conducted on a computer equipped with four GeForce RTX 2080 Ti graphics card and a pair of 14-core Intel E5-2690 v4 CPU.

### v4

@@ Appendix E Experimental details
Experimental setup.
We employed neural spline flows (Durkan et al.,, 2019) as the conditional density estimator, consisting of 8 layers. Following the flow structure settings in (Lueckmann et al.,, 2021), each layer consists of two residual blocks with 50 units and ReLU activation function, with 10 bins in each monotonic piecewise rational-quadratic transforms, and the tail bound was set to 5. Through empirical evaluation, we found that 8 layers provided sufficient flexibility for posterior estimation. Appendix E.4 compares 5-layer and 10-layer models, showing that 5 layers slightly degrade performance, while 10 layers perform similarly to 8 layers but with higher computational cost.
In the training process, the number of samples simulated in each round is $N=1000$ and there are $R=10$ rounds in total. In each round of training, we randomly select 5% of the newly generated samples as validation data. According to the early stop criterion proposed by Papamakarios et al., (2019), we stop training if the value of loss on validation data does not decrease after 20 epochs in a single round. For stochastic gradient descent optimizer, we choose the Adam (Kingma and Ba,, 2014) with the batchsize of 100, the learning rate of $5\times 10^{-4}$ and the weight decay is $10^{-4}$.
E.1 SIR model
E.1.1 Detailed results on synthetic data
In our experiments, we use the Gillespie algorithm (Gillespie,, 1977) to simulate the whole process over a duration of $T=160$ time units and record the populations at intervals of 1-time units. The prior distribution of $\beta$ is set to $\mathcal{N}(\log 0.4,0.5)$ and prior distribution of $\gamma$ is set to $\mathcal{N}(\log 0.125,0.2)$. The value of $K$ is configured as 1000000, while $N$ is set to 1000 for the number of observations.
To publish the privatized data about the infection process, we select the infectious group $I(t)$ at $L=10$ evenly-spaced points in time $[0,T]$, with the privacy parameters set to $n=1000$ and $m=1000$, which satisfied $\epsilon$-DP with $\epsilon=10$. The ground truth parameters are
@@ E.1.1 Detailed results on synthetic data
Figure 4 illustrates the convergence of the approximate posterior by SPPE and SPLE in each round, and compares our results with the SMC-ABC method, where we performed simulations up to $5\times 10^{5}$ times for the SMC-ABC method to generate the near exact ‘True’ posterior. The tolerance parameter $\epsilon_{t}$ in SMC-ABC was selected as $\epsilon_{t}=0.5\times 0.7^{t}$, where $t$ represents the iteration round, and the discrepancy function $\rho$ was chosen as the $\ell_{2}$ norm. Figure 5 depicts the results of the SMC-ABC method under the same performance metrics. Our method, after 10 rounds or $10^{4}$ simulations, achieved a similar performance as the SMC-ABC method with approximately $10^{5}$ simulations. The computational time costs comparison between our methods and SMC-ABC, as illustrated in Table 1, also reveals that our approaches require significantly fewer simulations, resulting in substantially lower simulation time expenditures compared to the SMC-ABC method. However, our methods require extra time for the training of normalizing flows, a duration dependent on the flow’s complexity. Overall, both SPPE and SPLE attain a low MMD more rapidly than SMC-ABC.
Table 1: Computational Cost to achieve MMD $<$ 0.1 in the SIR Model (Mean $\pm$ Standard Deviation)
Method Simulation Time (min) Network Training Time (min) Total Number of Simulations
@@
Figure 4: Detailed convergence of sequential posterior estimations given DP-protected infection trajectory under the SIR model. Each round entails $N=1000$ simulations. Orange: SPPE; red: SPLE; grey: SMC-ABC.
Figure 5: Approximation accuracy by SMC-ABC on the SIR model against the number of simulations.
Finally, we investigate the marginal posterior distributions $\pi(\beta\mid s_{\mathrm{dp}})$ and $\pi(\gamma\mid s_{\mathrm{dp}})$ in Figure 6, and conclude that SPPE and SPLE perform similarly well. In this example, the posterior approximated by SPPE is slightly more concentrated than those from SMC-ABC and SPLE.
Figure 6: Marginal posterior histograms of $\beta,\gamma$, and $\beta/\gamma$ in SIR model on synthetic data. Grey: SMC-ABC; orange: SPPE; red: SPLE. The vertical lines indicate true data generating parameters, set to mimic a measles outbreak.
Figure 7: Rate of convergence of MC and RQMC: The $x$-axis represents the number of samples $M$ used for integral estimation, and the $y$-axis shows the logarithmic value of the root-mean-square-error (RMSE). The results indicate that the RMSE of the MC method is approximately $\mathcal{O}(M^{-1/2})$, while the RMSE of the RQMC method is approximately $\mathcal{O}(M^{-1})$.
Finally, we evaluate the impact of the privacy loss budget on the performance of SPPE and SPLE. By setting the parameter $m$ to either $10^{4}$ or $10^{5}$, we can adjust the privacy loss budget to $\epsilon=1$ or $\epsilon=0.1$, respectively. Figure 8 illustrates the posterior approximation accuracy under different $\epsilon$ values. SPPE demonstrates rapid and stable convergence, while SPLE requires approximately 5 training rounds to converge.
A
B
Figure 8: Approximation accuracy by SPPE (orange) and SPLE (red) against the number of rounds with A. Privacy loss budget $\epsilon=1$. B. Privacy loss budget $\epsilon=0.1$. the error bars represent the mean with the upper and lower quartiles.
E.1.2 Inference on real disease outbreaks
We applied our inference methods (SPPE and SPLE) to analyze several real infectious disease outbreaks, namely influenza, Ebola, and COVID-19. For these experiments, we chose the privacy parameters $M=100$, $N=100$, and $L=10$, which result in an overall privacy loss budget $\epsilon=10$ according to Proposition 4.
influenza outbreak.
We utilized the dataset from a boarding school, obtained from https://search.r-project.org/CRAN/refmans/epimdr/html/flu.html. The total population in the school was $K=763$. We considered a daily time interval, and the observed private statistic $s_{\mathrm{dp}}^{o}$ is
@@ E.1.2 Inference on real disease outbreaks
Ebola outbreak in West Africa, 2014.
We analyzed datasets from three regions: a) Guinea, b) Liberia, and c) Sierra Leone. The dataset source is from https://apps.who.int/gho/data/node.ebola-sitrep. We assumed potential contact individuals of $K=100,000$. We selected 9 equally spaced time intervals of 120 days starting from 03/31/2014. The resulting observed private statistic $s_{\mathrm{dp}}^{o}$ is as follows
@@
COVID-19.
We examined the COVID-19 dataset for Clark County, Nevada. See https://usafacts.org/visualizations/coronavirus-covid-19-spread-map/state/nevada/county/clark-county/. We assumed a potential contact population of $K=100,000$. We selected 9 equally spaced time intervals of 24 days, starting from 09/07/2020. To obtain the number of currently infected individuals, we calculated the difference in the total confirmed cases with a time interval of 14 days from the original dataset. The resulting observed private statistic $s_{\mathrm{dp}}^{o}$ is:
@@
E.2 Bayesian linear regression
Data generating parameters.
Following the parameters settings in Ju et al., (2022), we model the predictors with $x_{0,i}\sim\mathcal{N}_{p}(m,\Sigma)$, where $\Sigma=I_{n}$ and $m=(0.9,\ -1.17)$. The outcomes $y$ satisfy $y\mid x_{0}\sim\mathcal{N}_{n}(x\beta,\sigma^{2}I_{n})$ where $\sigma^{2}=2$, and the prior for $\beta$ is independent normal $\mathcal{N}(0,1)$. We set the privacy loss budget $\epsilon=10$ and the number of subjects $n=100$. The ground truth parameters are denoted as
@@ E.2 Bayesian linear regression
Experimental results.
In Figure 9, we present a comparison of the performance of the SPPE and SPLE methods across different metrics as the number of simulation rounds increases. The SPLE method demonstrates a faster convergence in this task. Figure 10 displays the posterior after 10 rounds, where both the SPPE and SPLE methods achieve results that are close to the near exact posterior obtained by the SMC-ABC method.
Figure 9: Approximation accuracy by SPPE (orange) and SPLE (red) on the Bayesian linear regression model against number of rounds, the error bars represent the mean with the upper and lower quartiles.
Figure 10: Posterior comparison on the Bayesian linear regression model. Grey: SMC-ABC; orange: SPPE; red: SPLE. The vertical lines and black dots indicate true data generating parameters.
To further characterize the privacy-utility trade-off, we compare the private data posterior distributions under several levels of privacy loss budget, in Figure 11. The underlying confidential data $x$ is the same for each $\epsilon=0.1,0.3,1,3,10,30,100.$ For each privacy loss level, we simulate one private summary $s_{\mathrm{dp}}(\epsilon)$, and use SPPE and SPLE to approximate the private data posterior $\pi(\theta\mid s_{\mathrm{dp}};\epsilon).$ The two methods yield very similar results. More interestingly, we use the same confidential data $x$ as Ju et al., (2022) and the posterior distributions in Figure 11 follow a similar trend with that in Figure 2 of Ju et al., (2022). For larger privacy loss budget (smaller noise), confidential data are mainly corrupted by clamping, and the proposed methods can elevate the effect of this censoring bias, as $\mathbb{E}(\theta\mid s_{\mathrm{dp}})$ is closer to the confidential data expectation $\mathbb{E}(\theta\mid x).$ As $\epsilon$ gets closer to 0 (near perfect privacy), the privacy mechanism has injected so much noise into $s_{\mathrm{dp}}$ that the posterior distribution is more dispersed, and little information about $\theta$ can be learned based on observing $s_{\mathrm{dp}}.$
Finally, we compare the posterior means and 95% credible intervals obtained by different methods under various privacy loss budgets, as shown in Tables 23, and 4. When larger noise is added, the posterior from all methods converge toward the prior of the parameters. Additionally, the naive posterior approximation fails to compute results due to the non-positive definiteness of the covariance matrix. The corresponding Kolmogorov-Smirnov test results, presented in Figure 12, shows that RNPE exhibits a larger bias compared to other methods, while the results from DA-MCMC, SMC-ABC, SPPE, SPLE, and Gibbs-SS are closely aligned.
Table 2: Estimated posterior mean and 95% credible intervals for the linear regression example using various methods. Here privacy loss budget is set to $\epsilon=10$.
@@
DA-MCMC -0.62 (-2.50, 0.99) -2.72 (-3.74, -0.96) 0.54 (-1.06, 2.46)
SMC-ABC -0.59 (-2.28, 0.93) -2.44 (-3.67, -0.38) 0.85 (-0.88, 2.77)
SPPE -0.51 (-2.25, 1.07) -2.40 (-3.61, -0.30) 0.90 (-0.89, 2.85)
SPLE -0.64 (-2.31, 0.96) -2.61 (-3.65, -0.98) 0.64 (-0.93, 2.48)
Gibbs-SS -0.46 (-2.22, 1.39) -0.50 (-2.08, 1.28) 0.41 (-1.68, 2.13)
RNPE -1.40 (-3.59, 0.94) -2.15 (-3.34, 0.51) 0.29 (-2.11, 2.97)
Table 3: Estimated posterior mean and 95% credible intervals for the linear regression example using various methods. Here privacy loss budget is set to $\epsilon=1$.
$\beta_{0}$ $\beta_{1}$ $\beta_{2}$
Naive posterior approximation - - -
DA-MCMC -1.00 (-2.75, 0.85) -0.96 (-2.66, 0.89) 1.23 (-0.76, 2.88)
SMC-ABC -0.80 (-2.74, 1.17) -0.83 (-2.64, 1.23) 0.90 (-1.00, 2.58)
SPPE -0.93 (-2.80, 1.04) -0.90 (-2.62, 0.97) 1.11 (-0.71, 2.72)
SPLE -0.89 (-2.69, 1.06) -1.04 (-2.82, 0.91) 1.25 (-0.55, 2.88)
Gibbs-SS -0.92 (-2.72, 0.94) -0.67 (-2.21, 1.04) 1.04 (-0.62, 2.47)
RNPE -1.37 (-3.81, 2.07) -0.51 (-3.01, 2.53) 1.34 (-1.77, 3.76)
Table 4: Estimated posterior mean and 95% credible intervals for the linear regression example using various methods. Here privacy loss budget is set to $\epsilon=0.1$.
$\beta_{0}$ $\beta_{1}$ $\beta_{2}$
Naive posterior approximation - - -
DA-MCMC -0.12 (-1.98, 1.77) -0.13 (-2.13, 1.88) 0.13 (-1.91, 2.12)
SMC-ABC -0.09 (-1.97, 1.93) -0.10 (-2.09, 1.95) 0.11 (-1.90, 2.22)
SPPE -0.12 (-2.13, 1.73) -0.12 (-2.11, 1.83) 0.14 (-1.81, 2.09)
SPLE -0.04 (-2.01, 1.99) -0.06 (-2.00, 1.85) 0.12 (-1.86, 2.09)
Gibbs-SS -0.05 (-1.94, 1.95) -0.01 (-1.95, 1.97) 0.21 (-1.62, 1.89)
RNPE -0.22 (-3.56, 3.16) 0.13 (-3.12, 3.51) 0.34 (-3.12, 3.61)
A
B
Figure 12: Kolmogorov-Smirnov test statistics between approximations of posterior marginals. A. Privacy loss budget $\epsilon=1$. B. Privacy loss budget $\epsilon=0.1$.
E.3 Naïve Bayes log-linear model
Model description.
@@ E.3 Naïve Bayes log-linear model
The naïve Bayes log-linear model is a commonly used approach for modeling categorical data (Karwa et al.,, 2015). The input feature-vector, denoted as $x=(x_{1},\cdots,x_{K})$, consists of $K$ features, each taking values in the range $\{1,2,\cdots,J_{k}\}$. The output class, denoted as $y$, represents the target category and takes values in $\{1,2,\cdots,I\}$. The model assumes that the conditional probability of the input given the output, denoted as $p(x\mid y)$, can be factorized as the product of individual feature probabilities: $p(x\mid y)=\prod_{k=1}^{K}p(x_{k}\mid y)$. The model parameters are $p_{ij}^{k}$, which represents the probability $p(x_{k}=j\mid y=i)$, with prior $(p_{i,1}^{k},\cdots,p_{i,J_{k}}^{k})\sim\mathrm{Dirichlet}(\alpha_{i,1}^{k},\cdots,\alpha_{i,J_{k}}^{k})$ for all $i$ and $k$; and $p_{i}=p(y=i)$, with prior $(p_{1},\cdots,p_{I})\sim\mathrm{Dirichlet}(\alpha_{1},\cdots,\alpha_{I})$.
@@
We assume that $(n_{i,1}^{k},\cdots,n_{i,J_{k}}^{k})\sim\mathrm{Multinomial}(n_{i};p_{i,1}^{k},\cdots,p_{i,J_{k}}^{k})$ for all $i$ and $k$ and $(n_{1},\cdots,n_{I})\sim\mathrm{Multinomial}(n;p_{1},\cdots,p_{I})$. Here $n_{i,j}^{k}$ represents the counts #$\#(y=i,x_{k}=j)$. One sufficient statistics of the model is the proportion of the counts $r_{i,j}^{k}:=\frac{1}{n}n_{i,j}^{k}$, where $n=\sum_{i=1}^{I}\sum_{j=1}^{J_{k}}n_{i,j}^{k}$ for all $k$. To protect the privacy of the dataset, Laplace noise $e_{i,j}^{k}$ is added to the proportion of the counts, resulting in the privatized proportion $m_{i,j}^{k}=r_{i,j}^{k}+e_{i,j}^{k}$. When $e_{i,j}^{k}\sim\mathrm{Laplace}(0,\frac{2K}{n\epsilon})$, the released private statistic $\{m_{i,j}^{k}\}_{i,j,k}$ satisfied $\epsilon$-DP.
@@
Experimental results.
Figure 13 illustrates the performance of the SPPE and SPLE methods across four different metrics, and Figure 14 shows the marginal posterior histograms after 10 rounds, our methods stabilize in performance after round 3.
Figure 13: Approximation accuracy by SPPE (orange) and SPLE (red) on the log-linear model against the number of rounds, the error bars represent the mean with the upper and lower quartiles.
Figure 14: Marginal posterior histograms of the log-linear model. Grey: SMC-ABC; orange: SPPE; red: SPLE. The vertical lines indicate true data generating parameters.
E.4 Neural spline flows with different layers
To evaluate the impact of varying the number of layers in the neural spline flow (NSF) model, we conducted additional experiments using configurations with 5 and 10 layers, in addition to the existing 8-layer setup.
The results are shown in Figure 15. For the Bayesian linear regression task, the posterior approximation performance with 5-layer and 10-layer configurations is comparable to the 8-layer NSF. This indicates that a 5-layer NSF possesses sufficient flexibility to represent the posterior or likelihood. However, for the SIR model, the SPLE method with the 5-layer configuration converges slightly slowly compared to the 8-layer NSF, indicating that fewer layers may marginally affect convergence in some cases. Besides, as analyzed in Appendix D, increasing the number of layers also results in a linear increase in computational cost.
A
B
C
D
Figure 15: Comparison of approximation performance for NSF models with different layer configurations by SPPE (orange) and SPLE (red). A. 5-layer NSF on SIR model. B. 10-layer NSF on SIR model. C. 5-layer NSF on Bayesian linear regression model. D. 10-layer NSF on Bayesian linear regression model.
Appendix F Statement on Computing Resources
Our numerical experiments were conducted on a computer equipped with four GeForce RTX 2080 Ti graphics cards and a pair of 14-core Intel E5-2690 v4 CPUs.